OpenStudy (anonymous):

simplify: 3x-4y+3z-3x-y+2z

HanAkoSolo (jamierox4ev3r):

you add the like terms together

OpenStudy (anonymous):

combine like terms...xs with xs , ys with ys, etc

HanAkoSolo (jamierox4ev3r):

-4y and -y are like terms because of their variables

HanAkoSolo (jamierox4ev3r):

so, you would get -5y+5z

OpenStudy (anonymous):

6x-5y+z right?

HanAkoSolo (jamierox4ev3r):

the x-variables cancel each other out, and the z variables add up because they are both positive. It helps to break down the problem

OpenStudy (anonymous):

*5z sorry

HanAkoSolo (jamierox4ev3r):

3x-3x=0

OpenStudy (anonymous):

First, simplify the x variables 3x - 3x = 0 so -4y + 3z - y + 2z Next, simplify the y variables -4y - y = -5y so -5y + 3z + 2z Finally, simplify the z variables 3z + 2z And you'll get -5y + 5z

HanAkoSolo (jamierox4ev3r):

@ashleyjade do you understand?

OpenStudy (anonymous):

yes thank you!

OpenStudy (anonymous):

can you help with one more?